OpenStudy (anonymous):

How many times larger is the new volume of a cylinder if only the height tripled, and the radius remained the same

OpenStudy (lgbasallote):

\[V = \pi r^2 h\] if height will be tripled and radius stays the same \[\Large V_{new} = \pi r^2 (3h)\] \[\Large V_{new} = 3(\pi r^2 h)\] \[\Large V_{new} = 3(V_{old})\] does that help?

OpenStudy (goformit100):

pi r^2 (3h)

OpenStudy (anonymous):

so it triples ?

OpenStudy (lgbasallote):

is that how you understood what i said?

OpenStudy (lgbasallote):

and yes triples lol